OpenStudy (anonymous):

B is between A and C. Find BC if AB = 3, BC = 4x + 1, AC = 8. 1 5 -5 -1

OpenStudy (anonymous):

AC = AB + BC, so 8 = 3 +4x +1 4x = 4, so x = +1